# ReDDiT: Rehashing Noise for Discrete Visual Generation

### v2

@@ 4.1 Implementation
Pre-processing.
Following the setting in LlamaGen [39], we apply the ten-crop augmentation on images, and use pre-trained tokenizers to convert them to discrete tokens. We pick IBQ-f16 [35] tokenizer as default for its scalable and promising performance in generation tasks, which uses a $16\times 16$ downsampling ratio and converts a $256\times 256$ image into 256 discrete tokens. The tokenizer has a codebook with 16384 entries. The LlamaGen-f16 (used in Tab. 3) and LlamaGen-f8 tokenizer [39] (used in Tab. 2) are also used for comparison with recent discrete generation methods. All tokenizers are used out-of-the-box without any tuning or modification.
Training with Representation Alignment.
Recent study [46] has shown that the alignment of intermediate representations between diffusion transformers and vision encoders accelerates training convergence of diffusion models. Accordingly, the alignment is designed as a regularization term with $\lambda=0.5$. We extract diffusion transformer’s n-th layer intermediate feature $\textbf{h}^{[n]}(x_{t})$ and align it with the original image’s dinov2-b [25] encoded features $f_{\text{enc}}(x_{0}^{\text{ori}})$. The intermediate features are projected by a small trainable MLP $h_{\varphi}$. The $\text{sim}(\cdot,\cdot)$ computes the mean of element-wise cosine similarity between embeddings, as
This alignment was proposed for continuous diffusion models, and we firstly validate that it’s also suitable for training discrete models in Tab. 1 (a).
Model Training.
The proposed model is based on DiT [28] architecture, with reference to its discrete prediction version [31]. 2D-RoPE [38] and min-SNR [47] are applied for training efficiency. The model is optimized using the AdamW optimizer with a cosine decay. Training is conducted for 500k iterations on 8 NVIDIA H100 GPUs with a global batch size 1024. Class-conditional training is enabled using class embeddings and a drop-rate of 0.1 for generation with classifier-free guidance.
4.2 Determining Noise Capacity
Figure 4: Comparison of different noise capacities. We re-implemented training with MVTM and ReDDiT design with the same training recipe (LlamaGen-f16 as visual tokenizer, with codebook size $16384$). The generation quality and representation alignment trends are visualized.
The reformulated discrete diffusion dynamics defines transitioning from $\mathcal{E}_{d}$ to $\mathcal{M}_{m}$. Under this setting, it is necessary to empirically determine the optimal value of $m$ for a fixed tokenizer with vocabulary size $d$, as the latent representations learned by VAEs are variant. We keep the training setup fixed and conduct experiments w.r.t. the noise capacity $m$. We also visualize $\mathcal{L}_{\text{RepA}}$, which captures the degree of representation alignment [46] within the transformer.
4.3 Ablation Study
(a) (a) RepA choices
Layer RepA gFID
- $\times$ 7.24
4 5.72
8 5.41
12 5.82
16 6.12
(b) (b) General model design
Configuration gFID Prec. Rec.
MVTM (19) + RepA 6.83 0.75 0.39
    objective (20) 6.25 0.76 0.43
+ 2D-RoPE + min-SNR 5.97 0.78 0.44
+ rehash sampler 5.53 0.81 0.51
+ DFM refine 5.40 0.81 0.52
(c) (c) Sampling timeline
Steps Timeline gFID
20 linear 7.18
32 linear 6.43
20 arccos 5.04
20 square 7.39
20 cosine 4.91
Unless specified, all the models are trained on ImageNet $256\times 256$ under the default settings for 100k iterations for fair comparison. We use a constant guidance scale of 2.0 and 20 steps for generation, and report gFID $\downarrow$ and IS $\uparrow$ computed on 50K samples. Precision (Prec.$\uparrow$) and Recall (Rec.$\uparrow$) are also reported in general design for direct diversity comparison.
General Design.
We ablate the general choices of ReDDiT, which starts with a re-trained MVTM baseline methods (with LlamaGen-f16 and RepA for faster convergence as default) in Tab. 1. We also tested RepA supervision on different layers of ReDDiT. The results are consistent with the original study [46], which indicates that adding RepA objective at the 8-th of 24 layers results in best convergence for L-model. The objective and modifications are conducted with re-training. We replace the MVTM sampler with the proposed rehash sampler and DFM refinement for further improvement.
Sampling Timeline.
Recovering complete information from noise remains critical to diffusion-based models [22, 42]. Recent work shows MVTM’s non-linear scheduler for training is less critical when using high-capacity tokenizers. Evidence of time-invariance in DDMs [31, 36] further supports decoupling training from sampling. In our experiments, a linear scheduler—with constant signal-to-noise ratio decay—yields optimal training dynamics. Among the timeline discretization tested, Fig. 5, the cosine schedule is employed for our ReDDiT model for its optimal performance in Tab. 1.
4.4 Performance Comparison
We compare the proposed ReDDiT model with other generative models on the ImageNet-1K $256\times 256$ in Tab. 2. The IBQ tokenizer is used for the default L and XL models. We also utilize LlamaGen-f8 with 128 noise capacity to evaluate its high-resolution potentials (noted as ReDDiT-XLf8). We use a linear increasing guidance following the common practice of  [10].
Table 2: Performance comparison on class-conditional ImageNet 256$\times$256. Look-up free quantizers are beyond the scope of this paper. ft. indicates that the decoder is fine-tuned for quantized latents. Wall-clock inference time relative to ReDDiT-XL is reported.
Type Model Tokenizer Generator
@@ 4.4 Performance Comparison
Diff. LDM-4 [29] 4096×3 - 3.60 247.7 400M 250
DiT-XL/2 [28] 1024×4 - 2.27 278.2 675M 250 18
MDTv2 [10] 1024×4 - 1.58 314.7 676M 250 >18
MAR-H [20] 256×16 - 1.55 303.7 943M 256 10
MVTM MaskGIT [4] 256 1024 6.18 182.1 227M 8 0.2
TiTok-S-128ft. [45] 128 4096 1.97 281.8 287M 64 1.6
AR Taming-VQGAN [9] 256 1024 15.78 74.3 1.4B 256 8
RQ-Transformer [17] 256 16384 7.55 134.0 3.8B 64 8.5
ViT-VQGAN [43] 1024 8192 4.17 175.1 1.7B 1024 >10
LlamaGen-3B [39] 576 16384 2.18 263.3 3.1B 576 20
RandAR-XXL [27] 512 16384 2.15 322.0 1.4B 88 4
VAR-d30 [41] 680 4096 1.97 334.7 2.0B 10 0.5
DDM ITM [16] 1024 16384 5.30 183.0 546M 100 3
ReDDiT-L (ours) 256 16384 2.13 294.7 346M 20 0.5
ReDDiT-XL (ours) 256 16384 1.74 313.6 675M 32 1
ReDDiT-XLf8 (ours) 1024 16384 1.61 318.5 675M 64 2
Model VQ Tokenizer Info. Generator
@@
LlamaGen-LAR [39] LlamaGen-f16 [39] 2.19 8 343M 3.80
RandAR-LAR [27] 343M 2.55
OursDDM(ReDDiT-L) 346M 2.41
IBQ-BAR [35] IBQ-tokenizer [35] 1.37 256 343M 2.88
OursDDM(ReDDiT-L) 346M 2.13
Generation Quality.
As shown in Tab. 2, ReDDiT achieves the best performance among the compared discrete models. It outperforms the baseline (MaskGIT [4]) with significant margins (gFID: 2.13 vs 6.18 and IS: 294.7 vs. 182.1). It also outperforms the recent DDM method [16] and TiTok-S-128 [45], which is extensively fine-tuned on quantized latents. Compared with continuous diffusion models, ReDDiT exhibits higher efficiency and on-par performance, showing great potential for discrete generation. Note that the performance is achieved with a codebook size of 16384, validating ReDDiT’s effectiveness for large-vocabulary codebooks.
Efficiency.
ReDDiT is born with the high-efficiency advantage of discrete diffusion models. As shown in Tab. 2, the inference time of ReDDiT is slightly longer than MaskGIT, while the performance is overwhelming. With only 20 steps it achieves a competitive performance which diffusion models uses more than 250 steps to achieve. It also demonstrates comparable results with significantly smaller model size than recent VAR [41] method.
Besides the major comparison, we also conduct an experiment that utilizes the identical tokenizer in previous auto-regressive (AR) works and validate our method’s effectiveness. As can be seen in Tab. 3, ReDDiT outperforms AR methods in generation tasks across different tokenizers.
4.5 Qualitative Results
Figure 6 presents representative class-conditional samples generated by the proposed ReDDiT model. The outputs across diverse image classes consistently exhibit both high fidelity and diversity. We further demonstrate ReDDiT’s editing capability in Figure 7, highlighting its bi-directional perceptual competence. Following MaskGIT [4], we replace a region of the input image with noise tokens and employ the same generation pipeline to inpaint the missing content, conditioned on a class label $c$.
@@ 4.5 Qualitative Results
5 Conclusion
We proposed ReDDiT, a discrete visual generative model built upon a principled discrete diffusion architecture with novel noise designs and efficient sampling strategies. Our key contribution lies in the integration of rehashing noise and samplers, which together ensure both diversity and low discrepancy throughout the generative process. By introducing rehashing noise, ReDDiT enriches the potential paths that latent variables can traverse during training, enhancing representational capacity. Extensive experiments demonstrate that discrete generative models can achieve performance on par with their continuous counterparts while offering significantly improved efficiency. This work paves a promising way for discrete generative modeling and offers fresh insights toward unifying visual and language generation—a path we leave for future exploration.

### v3

@@ Appendix C Discrete Diffusion with Rehashing Noise
Appendix D Sampling from Learned Networks
We present a detailed version of discrete flow matching (DFM) sampler 3 and ours 4, and discuss the integration of them. Fig. 7 presents a quantitative comparison of the vanilla DFM sampler, our proposed rehash sampler, and a hybrid strategy that combines both approaches by incorporating selected DFM steps into the rehash trajectory. All methods are evaluated using identical model weights, as the training objectives are compatible due to their shared time-weighted loss formulation.
The rehash sampler exhibits stronger overall performance than DFM, especially in the 15–32 step range, where it achieves low and stable gFID scores. This suggests that our modification enables more efficient decoding trajectories without sacrificing sample quality. The hybrid variant, which integrates only the middle and final steps of the DFM update into the rehash schedule, also delivers consistent gains over the vanilla DFM, suggesting that partial refinement from DFM is beneficial even when the majority of the trajectory is governed by our rehash dynamics.
By leveraging shared gradual decoding infrastructure, the hybrid approach enables practical integration of DFM refinement into the ReDDiT framework with minimal overhead. As noted in the main paper, this leads to a $\sim$0.1 improvement in gFID on ImageNet-1K, reinforcing the complementary strengths of the two samplers. We leave the comprehensive study on the optimal integration of different samplers for future exploration.
Figure 7: Generation quality comparison with DFM methods. The experiments are conducted on ReDDiT-L with a constant classifier-free guidance ($\text{cfg}=2.0$).
Algorithm 3 DFM Sampling Stepwise Pseudo Code
@@ Appendix D Sampling from Learned Networks
2:Compute jump probabilities: $j_{t}\leftarrow 1-\alpha_{t}$, $j_{s}\leftarrow 1-\alpha_{t-\Delta t}$
3:Determine guidance scale $\omega$ from schedule
4:Obtain logits $\text{logits}_{\text{cond}}$, $\text{logits}_{\text{uncond}}$ via forward pass
@@
[…]
7:Sample $\hat{x}_{0}\sim p_{x_{0}}$ using categorical sampling
8:Construct one-hot encodings: $\delta_{x_{0}},\delta_{x_{t}}$
[…]
11:Overwrite $u$ in masked range with corrective terms
12:Mask entries already present in $x_{t}$ from $u$
13:Compute total transition intensity: $\lambda\leftarrow\sum u$, elementwise
14:Draw Bernoulli mask: $M\sim\text{Bernoulli}(1-\exp(-\lambda))$
15:For each masked position in $M$, sample from categorical $u$ to obtain updated $x_{s}$
16:return $x_{s}$
Algorithm 4 Rehash Sampling (ours) Stepwise Pseudo Code
[…]
2:Indentify the masked tokens: $M\leftarrow[x_{t}\in\mathcal{M}_{m}]$
3:Rehash $x_{t}$: $x_{t}\leftarrow M\cdot\texttt{random\_shuffle}(\text{mask\_vocab})+(1-M)\cdot x_{t}$.
4:Compute move coefficients: $k_{t}\leftarrow 1-\alpha_{t}$, $k_{s}\leftarrow 1-\alpha_{t-\Delta t}$
5:Determine guidance scale $\omega$ from schedule
6:Obtain logits $\text{logits}_{\text{cond}}$, $\text{logits}_{\text{uncond}}$ via forward pass
[…]
9:Set mask probability: $p_{\text{mask}}\leftarrow k_{s}$
10:Construct proposal distribution: $q_{x_{s}}\leftarrow\frac{k_{t}-k_{s}}{k_{t}}\cdot p_{x_{0}}$
11:Overwrite mask token logits:
$q_{x_{s}}[:,:,m{:}]\leftarrow 0,\quad q_{x_{s}}[:,:,m]\leftarrow\frac{p_{\text{mask}}}{k_{t}}$
where $m$ is the start of mask token index
12:Sample $\hat{x}\sim q_{x_{s}}$ using categorical sampling
13:Identify preserved tokens: $c\leftarrow[x_{t}<m]$
14:Combine result: $x_{s}\leftarrow c\cdot x_{t}+(1-c)\cdot\hat{x}$
15:return $x_{s}$
Appendix E Experiment Details
We provide detailed training and generation configurations for ReDDiT in Table 4. Our method incorporates DINOv2-B for representation alignment, which requires computing image features during the forward pass (only activated during training). This introduces an overhead, making training roughly 1.2× slower than solely on discrete tokens. However, this additional cost is offset by faster convergence and improved stability, particularly in early training stages.
The use of quantized latents allows for larger batch sizes under limited GPU memory, making our approach more accessible for low-resource settings. Additionally, aligning discrete codes with semantic features improves the quality and diversity of learned representations. Overall, our design balances computational efficiency with model performance, making it a practical choice for both research and deployment.
Table 4: Experiment details for ReDDiT on ImageNet-1K. Vari. refers to a time-variant growing guidance scale following MDTv2, which is a common practice for diffusion models.
Setting ReDDiT-L (Ablation) ReDDiT-L ReDDiT-XL ReDDiT-XLf8
Hidden Size 1024 1024 1280 1280
Transformer Block 24 24 28 28
Attention Head 16 16 20 20
Image Tokenizer LlamaGen-f16 IBQ-f16 IBQ-f16 LlamaGen-f8
Codebook Size 16384 16384 16384 16384
Noise Capacity 128 1024 1024 128
Sequence Length 256 256 256 1024
RepA Latent Size 16×16 16×16 16×16 32×32
Batch Size 64 64 32 16
Global Batch Size 1024 1024 1024 1024
LR scheduler Cosine Decay Cosine Decay Cosine Decay Cosine Decay
Learning Rate 3e-4 3e-4 3e-4 4e-4
Minimal LR 1e-5 1e-5 1e-5 1e-5
Warmup Steps 2k 2k 2k 2k
Training Steps 500k 500k 500k 500k
Training Time $\sim$1 day $\sim$1 day $\sim$2 days $\sim$3 days
Generation CFG (Vari.) 1.0-5.0 1.0-6.5 1.0-6.5 1.0-5.5
Appendix F Accelerating ReDDiT
Recent efforts on scaling and accelerating discrete diffusion models are making this generative paradigm more practical than theoretical attempts. We adapt the dLLM-Cache Liu et al. (2025) design into our framework, which efficiently reuses intermediate computations without compromising model performance. Since the condition is modulated using AdaLN and introduces minimal calculation, we do not activate $K_{p}$ (cache for prompt). As the decoding of visual sequence varies with time more quickly than language decoding, we implement the cache for response with small values like $K_{r}=2\ \text{or}\ 4$, which means the $K$ and $V$ of transformer layer is updated every $2$ or $4$ decoding steps instead of per step. As shown in Tab. 5, the inference speed is boosted up to 2 times with minimal performance drop, which makes our largest model ReDDiT-XLf8 comparable to diffusion models with accelerated solvers.
Table 5: Acceleration of ReDDiT using response cache $K_{r}$.
Model Config Performance
[…]
ReDDiT-L 32 2 $\times 1.33$ 2.28 ($\Delta=0.15$)
ReDDiT-XL $\times 1.52$ 1.88 ($\Delta=0.14$)
ReDDiT-XL 64 4 $\times 2.17$ 1.83 ($\Delta=0.09$)
ReDDiT-XLf8 $\times 2.56$ 1.71 ($\Delta=0.10$)
Appendix G Qualitative Results
We provide a comparison between MVTM and our method on generated images, and more samples of ReDDiT’s generation in Fig. 8.